OpenStudy (anonymous):

Solve the equation. Show your work. 45 = 3b + 69

OpenStudy (anonymous):

isolate 3b first.

OpenStudy (anonymous):

thus, get rid of 69 on right side.

OpenStudy (anonymous):

69 is being added on right side, thus, we will use opposite operation of addition,

mathslover (mathslover):

@violy meant to say that : \[\large{\color{red}{45-69=3b+69-69}}\] what do you get @daboss33 ?

OpenStudy (anonymous):

45 - 69 = -24

mathslover (mathslover):

where did 3b go @daboss33 ?

OpenStudy (anonymous):

i think daboss is doing left side first.

OpenStudy (anonymous):

you put it aside til you get the first answer then you bring it back??

OpenStudy (anonymous):

do you mean, -24=3b ?

OpenStudy (anonymous):

<-- scracthing head umm no

OpenStudy (anonymous):

<-- lost again

OpenStudy (anonymous):

let me give you an example: 6 = 2b + 4 . First thing is to subtract 4 from both sides. thus, 6-4 = 2b + 4 -4 ===> 2 = 2b + 0 or 2 = 2b.

mathslover (mathslover):

ok leave that @daboss33 if you have 69 cakes and you ate all the cakes , what you have now?

OpenStudy (anonymous):

lol a fat stomach and nothing left to snack on

OpenStudy (anonymous):

also note that, any number minus itself is zero.

mathslover (mathslover):

@daboss33 so no cakes left right?

OpenStudy (anonymous):

ok i understand that much its the letters that confuse me

OpenStudy (anonymous):

right

mathslover (mathslover):

so wht we had : \[\large{\color{blue}{45-69 = 3b + 69-69}}\] \[\large{45-69=3b+0}\] right ?

OpenStudy (anonymous):

ok i think i understand it but i still have to find out what 3b is right?

mathslover (mathslover):

right so first tell me 45 - 69 = ?

OpenStudy (anonymous):

3b

mathslover (mathslover):

no calculate that

OpenStudy (anonymous):

-24

mathslover (mathslover):

great

mathslover (mathslover):

now we have \[\large{\color{green}{-24=3b}}\] divide both sides by 3

mathslover (mathslover):

that is : \[\large{\color{blue}{\frac{-24}{3}=\frac{3b}{3}}}\]

mathslover (mathslover):

what do you get now?

OpenStudy (anonymous):

ok i got stuck there i know 3 go into 3 1 time but you cant divide a neg number with 3

mathslover (mathslover):

we can do that @daboss33

mathslover (mathslover):

can we write -24 as -1 * 24 ?

OpenStudy (anonymous):

so it would be -1/b then??

mathslover (mathslover):

\[\large{\color{brown}{\frac{-1*24}{3}=\frac{3b}{3}}}\] \[\large{\color{blue}{-1*\frac{24}{3}=\frac{3}{3}*b}}\] \[\large{\color{brown}{-1*8=b}}\] \[\large{\color{orange}{-8=b}}\]

mathslover (mathslover):

got it?

OpenStudy (usukidoll):